OpenStudy (anonymous):

Evaluate 2x − y when x = 3 and y = −2

OpenStudy (anonymous):

Just plug the coff. of the variable in the equation : 2(3)-(-2)= ?

OpenStudy (anonymous):

just plug in the values.

OpenStudy (anonymous):

so what the anwers

OpenStudy (anonymous):

Eyad gave you the problem, and if you know how to multiply, then you should be able to do it

OpenStudy (anonymous):

9

OpenStudy (anonymous):

try again...

OpenStudy (anonymous):

5

OpenStudy (anonymous):

-

OpenStudy (anonymous):

2(3) - (-2) = start with 2 x 3 = what ?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

what is 2 x 3 ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

let me show you... 2(3) - (-2) -- a negative times a negative becomes a positive 6 + 2 = your answer So what is 6 + 2 ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Correct !!! Good job !

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

how is that incorrect ? 2x - y when x = 3 and y = - 2 plug in 3 for x and -2 for y 2x - y 2(3) - (-2) 6 + 2 = 8 That should be correct.

OpenStudy (anonymous):

Did you write the problem correctly ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why did you put -5 ? Is that supposed to be the correct answer ?

OpenStudy (anonymous):

2x - y - Z 2(-2) - (3) - (-2) -4 - 3 + 2= -5

OpenStudy (anonymous):

Plug the numbers in

OpenStudy (anonymous):

you didn't write the problem correctly. You said y was -2 and there wasn't even a z in the problem.

OpenStudy (anonymous):

and x is supposed to be 3

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

next time make sure you write the problem down correctly and you will get the right answer

OpenStudy (anonymous):

im correct

OpenStudy (anonymous):

no you are not.....that is, if you are going by your original problem

OpenStudy (anonymous):

so what the answer then???

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

what is the correct problem including what the variables equal ?

OpenStudy (anonymous):

Is the problem 2x - y - z where x = 3, y = -2, and what does z equal ?

OpenStudy (anonymous):

or are you supposed to find z ?

OpenStudy (anonymous):

Evaluate 2x − y − z when x = −2, y = 3, and z = −2

OpenStudy (anonymous):

2x - y - Z 2(-2) - (3) - (-2) -4 - 3 + 2= (-5 answer)

OpenStudy (anonymous):

ok...let me do this... 2x - y - z 2(-2) - 3 - (-2) -4 - 3 + 2 - 7 + 2 - 5 You are correct...it is - 5. Next time, write the correct problem down and it would make this much easier

OpenStudy (anonymous):

@kelliegirl33 for your patience you deserve all the medals in the world.

OpenStudy (anonymous):

lol...thank you Eleven17....sometimes it is not so easy